Can I please have your assistance? If f(x) = 2x +1 and g(x) = x2, what is f o g? What is g o f?
Answer by Theo(13342)

fog(x) = f(g(x))
gof(x) = g(f(x))

you have:

f(x) = 2x+1
g(x) = x^2

fog(x) = f(g(x)) = f(x^2) = 2*(x^2) + 1. = 2x^2 + 1
you are simply replacing x in f(x) with g(x).
x is the argument in f(x).
g(x) is the argument in f(g(x))
since g(x) = x^2, then f(g(x)) becomes f(x^2) which becomes 2*(x^2) + 1 which becomes 2x^2 + 1.

gof(x) = g(f(x)) = g(2x+1) = (2x+1)^2 = 4x^2 + 4x + 1
you are simply replacing x in g(x) with f(x).
x is the argument in g(x).
f(x) is the argument in g(f(x))
since f(x) = 2x+1, then g(f(x)) becomes g(2x+1) which becomes (2x+1)^2 which becomes 4x^2 + 4x + 1